A steel thick-walled gas cylinder used for storing compressed air is fitted with a safety


valve which triggers at a pressure of 1.0 x 106 Pa. The cylinder contains air at 17°C and


0.8x106 Pa. At what temperature will the safety valve trigger?

1
Expert's answer
2021-12-31T08:55:05-0500

According to the combined gas law for constant volume:


"T_T=P_TT_0/P_0,\\
T_T=10^6·(273+18)/(0.8·10^6)=364\text{ K},"

or 364-273=91°C.